### v2

@@ 2.2 Minimal Pair Sets
We create the evaluation set using a pronunciation dictionary and speech synthesis. We first define a graph where the vertices of the graph are the words in the pronunciation dictionary and each edge is a phoneme swap that defines a minimal pair according to that pronunciation dictionary. Cliques (vertex subsets where every pair of vertices is adjacent) in this graph are minimal pair sets. We look at two accents of English, UK (Standard Southern British English) and US (General American), and one accent of Russian (Moscow). We choose Russian as it contains phonemically contrastive palatilization and a high central vowel that is not present in the accents of English we have chosen. We use the english_us_mfa, english_uk_mfa and russian_mfa pronunciation dictionaries from the Montreal Forced Aligner [9] to define the graph and the Bron-Kerbosch algorithm [10] for finding cliques. We then manually selected 10 minimal pair sets for each condition. 5 of these sets were vowel minimal sets designed to cover the basic monophthongs of each accent/language. The other 5 were consonant minimal sets, which contained a randomly selected version of each voiced/unvoiced consonant in the set. To synthesise the minimal pairs, we use Google’s Standard speech synthesis model  [^1] https://cloud.google.com/text-to-speech. We chose this model as it synthesises single words with minimal prosodic variation whilst retaining distinct speaker attributes. This set is used to evaluate the AAI models in Section 3.
@@ 3.1 Single and Multi-Speaker Models
The single and multi-speaker datasets used in this study are the dataset from [12] (henceforth called Medina) and XRMB. We use the whole dataset when training the single-speaker models and the train/validation split from [7] for training models on XRMB. We adapt WavLM Large in two ways, firstly with a linear layer and secondly with LoRA. Whilst there is some evidence that features from the intermediate layers of WavLM might perform better for inversion [13], we find that final-layer and intermediate layer features encode similar knowledge about the similarity between different consonants. We DTW aligned every combination of speaker (except the same speakers) and consonant from a subset of the XRMB data in Section 2.1 using a cosine distance cost function and present the average cosine similarities between DTW aligned features for each consonant combination in Figure 3. The articulatory representations are more similar for contrasts which are similar in terms of place and manner (e.g. bilabials), whereas 9th and final layer WavLM features are acoustically discriminant, with voiceless plosives /t/ and /k/ being more similar than the voicing contrast /k/ and /g/. We therefore opt to only use final layer features in the experiments as there is evidence that there is less speaker information in these layers [14].
@@ 5 Results
From the results given in Table 1, it is clear that the multi-speaker models outperformed the single-speaker models in terms of minimal pair classification and intelligibility, with linear-adapted models performing marginally better than LoRA-adapted models. All models achieved voicing scores that were much higher than the random WavLM projection benchmark of $\sim$50%, showing that all models learnt a form of articulatory representation, although none of the linear models achieved 100% voicing score. A crucial result was that models which had higher correlation on the XRMB set did not perform better on the classification and intelligibility tasks. Additionally, there was a strong negative correlation between the average minimal pair classification performance and WER, with correlations of -0.96 for the English and -0.89 for the Russian set. We found in general that models which scored highly on the minimal pair classification task would output clear articulatory targets, as in Figure 4.
@@
The target positions from linear models had noticeably higher variance than those from LoRA-adapted models. We haven’t examined the articulatory dynamics of these models in any detail, but the higher variance of the linear models is likely a result of the fact that the predicted trajectories are not smooth, as is shown in Figure 5. Although we use a low-pass filter in the minimal pair analysis, the lack of smoothness in the output trajectories may affect the final target positions. As is visible in Figure 5, the LoRA-adapted models do produce smooth trajectories, but whether they are valid articulatory dynamics [24] is a topic for future work. Once a method is created for evaluating dynamics, we could also use critical point analysis [8] to get the exact targets for classification.

### v3

@@ 2.2 Minimal Pair Sets
We create the evaluation set using a pronunciation dictionary and speech synthesis. We first define a graph where the vertices of the graph are the words in the pronunciation dictionary and each edge is a phoneme swap that defines a minimal pair according to that pronunciation dictionary. Cliques (vertex subsets where every pair of vertices is adjacent) in this graph are minimal pair sets. We look at two accents of English, UK (Standard Southern British English) and US (General American), and one accent of Russian (Moscow). We choose Russian as it contains phonemically contrastive palatilization and a high central vowel that is not present in the accents of English we have chosen. We use the english_us_mfa, english_uk_mfa and russian_mfa pronunciation dictionaries from the Montreal Forced Aligner [9] to define the graph and the Bron-Kerbosch algorithm [10] for finding cliques. We then manually selected 10 minimal pair sets for each condition. 5 of these sets were vowel minimal sets designed to cover the basic monophthongs of each accent/language. The other 5 were consonant minimal sets, which contained a randomly selected version of each voiced/unvoiced consonant in the set. To synthesise the minimal pairs, we use Google’s Standard speech synthesis model  [^1] https://cloud.google.com/text-to-speech. We choose this model as it synthesises single words with minimal prosodic variation whilst retaining distinct speaker attributes. This set is used to evaluate the AAI models in Section 3.
@@ 3.1 Single and Multi-Speaker Models
The single and multi-speaker datasets used in this study are the dataset from [12] (henceforth called Medina) and XRMB. We use the whole dataset when training the single-speaker models and the train/validation split from [7] for training models on XRMB. We adapt a WavLM Large model in two ways, firstly with a linear layer and secondly with LoRA. Whilst there is some evidence that features from the intermediate layers of WavLM might perform better for inversion [13], we find that final-layer and intermediate layer features encode similar knowledge about the similarity between different consonants. We DTW aligned every combination of speaker (except the same speakers) and consonant from a subset of the XRMB data in Section 2.1 using a cosine distance cost function and present the average cosine similarities between DTW aligned features for each consonant combination in Figure 3. The articulatory representations are more similar for contrasts which are similar in terms of place and manner (e.g. bilabials), whereas 9th and final layer WavLM features are acoustically discriminant, with voiceless plosives /t/ and /k/ being more similar than the voicing contrast /k/ and /g/. We therefore opt to only use final layer features in the experiments as there is evidence that there is less speaker information in these layers [14].
@@ 5 Results
From the results given in Table 1, it is clear that the multi-speaker models outperformed the single-speaker models in terms of minimal pair classification and intelligibility, with linear-adapted models performing marginally better than LoRA-adapted models. All models achieved voicing scores that were much higher than the random WavLM projection benchmark of $\sim$50%, showing that all models learnt a form of articulatory representation, although none of the linear models achieved 100% voicing score. A crucial result was that models which had higher correlation on the XRMB evaluation set did not perform better on the classification and intelligibility tasks. Additionally, there was a strong negative correlation between the average minimal pair classification performance and WER, with correlations of -0.96 for the English and -0.89 for the Russian set. We found in general that models which scored highly on the minimal pair classification task would output clear articulatory targets, as in Figure 4.
@@
The target positions from linear models had noticeably higher variance than those from LoRA-adapted models. We haven’t examined the articulatory dynamics of these models in any detail, but the higher variance of the linear models is likely a result of the fact that the predicted trajectories are not smooth, as is shown in Figure 5. Although we use a low-pass filter in the minimal pair analysis, the lack of smoothness in the output trajectories may affect the final target positions. As is also visible in Figure 5, the LoRA-adapted models do produce smooth trajectories, but whether they are valid articulatory dynamics [24] is a topic for future work. Once a method is created for evaluating dynamics, we could also use critical point analysis [8] to get the exact targets for classification.